OpenStudy (anonymous):

the length of a rectangular room is six times as long as the width.the perimeter is 84 yards find the width

OpenStudy (saifoo.khan):

width = x, length = 6x Perimeter = 2(L+B) = 2(6x + x)

OpenStudy (saifoo.khan):

width = x, length = 6x Perimeter = 2(L+B) = 2(6x + x) You know the perimeter already. So plug it in.

OpenStudy (anonymous):

okay so its 84=2(6x + x) now to i solve

OpenStudy (saifoo.khan):

Solve for x.

OpenStudy (anonymous):

x=6 is this correct? and what else do i have to do

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

Right. Now that is x. And x was width. So your width is 6.

OpenStudy (anonymous):

oh okay thank you very helpful just wanted to make sure the answer was correct

OpenStudy (saifoo.khan):

Now, you can find the length too, right?

OpenStudy (anonymous):

yea which is 6x if x=6 replace it meaning 6(6)= 36

OpenStudy (saifoo.khan):

Perfect.